OpenStudy (anonymous):

Please help! Given that f(x) = x2 + 3x + 6 and g(x) = the quantity of x plus two, over three, solve for f(g(x)) when x = 1.

OpenStudy (anonymous):

−1 0 3 4 answer choices

OpenStudy (campbell_st):

well 1st find g(1) by substituting x = 1 into g(x) then when you get that value substitute it into f(x)

OpenStudy (campbell_st):

but the answer isn't on the list

OpenStudy (anonymous):

so like f(X)=1+2/3

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well I've assumed \[g(x) = \frac{x + 2}{3}\] so what do you get when x = 1

OpenStudy (anonymous):

its actually x-3 over 2

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well that changes it so \[g(x) = \frac{x - 3}{2}\] substitute x = 1 and what do you get..?

OpenStudy (campbell_st):

then substitute the answer into f(x)